# Exact Exploratory Bi-factor Analysis: A Constraint-based Optimisation Approach

### v1

@@ C.2 Identifiability of Estimated Bi-factor Structure in Real Data Example
Thirdly, we denote by $\mathcal{B}_{5}^{*}=\{i_{5},i_{6}\}$, $\mathcal{B}_{6}^{*}=\{i_{7},i_{8}\}$ and $\mathcal{B}_{7}^{*}=\{i_{9},i_{10}\}$. Since $\Lambda$, $\Phi$ and $\Psi$ satisfy Condition 4, there exists three types of possible of bi-factor structure of $\mathcal{B}_{5}$, $\mathcal{B}_{6}$ and $\mathcal{B}_{7}$: (1)$\mathcal{B}_{5}=\{i_{5},i_{6}\}$, $\mathcal{B}_{6}=\{i_{7},i_{8}\}$ and $\mathcal{B}_{7}=\{i_{9},i_{10}\}$; (2) $\mathcal{B}_{5}=\{i_{5},i_{10}\}$, $\mathcal{B}_{6}=\{i_{6},i_{7}\}$ and $\mathcal{B}_{7}=\{i_{8},i_{9}\}$; (3) $\mathcal{B}_{5}=\{i_{5},i_{6}\}$, $\mathcal{B}_{6}=\{i_{7},i_{9}\}$ and $\mathcal{B}_{7}=\{i_{8},i_{10}\}$.
We now show that the second case is not allowed. We consider the off diagonal part of $\mathcal{B}_{1}^{*}$, $\mathcal{B}_{5}^{*}$, $\mathcal{B}_{6}^{*}$ and $\mathcal{B}_{7}^{*}$ rows and columns of $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}$ and $\Lambda\Phi(\Lambda)^{\top}+\Psi$. Firstly, we have $\Lambda^{*}[\mathcal{B}^{*}_{1},\{1\}]\lambda^{*}_{i_{5},1}+\phi^{*}_{2,6}\Lambda^{*}[\mathcal{B}^{*}_{1},\{2\}]\lambda^{*}_{i_{5},6}=\Lambda[\mathcal{B}^{*}_{1},\{1\}]\lambda_{i_{5},1}+\phi_{2,6}\Lambda[\mathcal{B}^{*}_{1},\{2\}]\lambda_{i_{5},6}$. Since $\Lambda[\mathcal{B}^{*}_{1},\{1,2\}]$ are linear independent and we have proved $\Lambda[\mathcal{B}^{*}_{1},\{1\}]=\Lambda[\mathcal{B}_{1},\{1\}]$ and $\Lambda[\mathcal{B}^{*}_{1},\{2\}]=\Lambda[\mathcal{B}_{1},\{2\}]$, we have $\lambda^{*}_{i_{5},1}=\lambda_{i_{5},1}$. Similarly, we have $\lambda^{*}_{i_{k},1}=\lambda_{i_{k},1}$ for $k=5,\ldots,10$. Next, we have $\lambda^{*}_{i_{5},1}\lambda^{*}_{i_{6},1}+\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{6},6}=\lambda_{i_{5},1}\lambda_{i_{6},1}+\phi_{6,7}\lambda_{i_{5},6}\lambda_{i_{6},7}$, which leads to $\phi_{6,7}=\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{6},6}/(\lambda_{i_{5},6}\lambda_{i_{6},7})$. We can similarly have $\phi_{67}=\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{6},6}/(\lambda_{i_{5},6}\lambda_{i_{6},7})=\phi^{*}_{6,7}\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{7},7}/(\lambda_{i_{5},6}\lambda_{i_{7},7})=\phi^{*}_{6,8}\lambda^{*}_{i_{6},6}\lambda^{*}_{i_{10},8}/(\lambda_{i_{6},7}\lambda_{i_{10},6})=\phi^{*}_{7,8}\lambda^{*}_{i_{7},7}\lambda^{*}_{i_{10},8}/(\lambda_{i_{7},7}\lambda_{i_{10},6})$. Combing these equations, we have $\phi^{*}_{6,7}\phi^{*}_{6,8}=\phi^{*}_{7,8}$. Also, by Condition 4, these equations lead to $\phi^{*}_{6,7},\phi^{*}_{6,8},\phi^{*}_{7,8}\neq 0$. In the same manner, we have $\phi^{*}_{6,8}\phi^{*}_{7,8}=\phi^{*}_{6,7}$ and $\phi^{*}_{6,7}\phi^{*}_{7,8}=\phi^{*}_{6,8}$. Then, we have $|\phi^{*}_{6,7}|=|\phi^{*}_{6,8}|=|\phi^{*}_{7,8}|=1$, which violates the assumption that $\Phi^{*}$ is positive definite.
Next, we show that the third case is not allowed. Similar to the proof in the second case, we first have $\lambda^{*}_{i_{k},1}=\lambda_{i_{k},1}$ for $k=5,\ldots,10$. Then we have the following 6 equations:
$\lambda^{*}_{i_{7},7}\lambda^{*}_{i_{8},7}=\phi_{7,8}\lambda_{i_{7},7}\lambda_{i_{8},8},$ (C.9)
$\phi^{*}_{7,8}\lambda^{*}_{i_{7},7}\lambda^{*}_{i_{9},8}=\lambda_{i_{7},7}\lambda_{i_{9},7},$
$\phi^{*}_{7,8}\lambda^{*}_{i_{7},7}\lambda^{*}_{i_{10},8}=\phi_{7,8}\lambda_{i_{7},7}\lambda_{i_{10},8},$
$\phi^{*}_{7,8}\lambda^{*}_{i_{8},7}\lambda^{*}_{i_{9},8}=\phi_{7,8}\lambda_{i_{8},8}\lambda_{i_{9},7},$
$\phi^{*}_{7,8}\lambda^{*}_{i_{8},7}\lambda^{*}_{i_{10},8}=\lambda_{i_{8},8}\lambda_{i_{10},8},$
$\lambda^{*}_{i_{9},8}\lambda^{*}_{i_{10},8}=\phi_{7,8}\lambda_{i_{9},7}\lambda_{i_{10},8}.$
@@
With Condition 4, according to the first equation above, we have $\phi_{7,8}\neq 0$. By the 6 equations, we also have $(\phi^{*}_{7,8})^{4}\phi_{7,8}^{4}=\phi_{7,8}^{2}$. Then we have $|\phi^{*}_{7,8}|=|\phi_{7,8}|=1$, which violates the assumption that $\Phi^{*}$ is positive definite.
Thus, only the case such that $\mathcal{B}_{5}=\{i_{5},i_{6}\}$, $\mathcal{B}_{6}=\{i_{7},i_{8}\}$ and $\mathcal{B}_{7}=\{i_{9},i_{10}\}$ is possible. Similar to the proof before, we have $\lambda^{*}_{i_{k},1}=\lambda_{i_{k},1}$ for $k=5,\ldots,10$. Similar to the proof of Proposition 1 of Fang et al., (2021), by Condition 4, there exists some $g\in\{1,\ldots,7\}$ such that $\phi_{6,1+g}\neq 0$, we then have $c(\lambda^{*}_{i_{5},6},\lambda^{*}_{i_{6},6})=(\lambda_{i_{5},6},\lambda_{i_{6},6})$ for some constant $c$. Since $\lambda^{*}_{i_{5},6}\lambda^{*}_{i_{6},6}=\lambda_{i_{5},6}\lambda_{i_{6},6}$, we have $c=1$ or $-1$, which leads to $(\lambda^{*}_{i_{5},6},\lambda^{*}_{i_{6},6})=(\lambda_{i_{5},6},\lambda_{i_{6},6})$ or $(\lambda^{*}_{i_{5},6},\lambda^{*}_{i_{6},6})=-(\lambda_{i_{5},6},\lambda_{i_{6},6})$. Then, we further have $\Psi_{i_{5}}=\Psi^{*}_{i_{5}}$ and $\Psi_{i_{6}}=\Psi^{*}_{i_{6}}$. The proof for observations $i_{7},\ldots,i_{10}$ is similar. Then, the proof is completed.

### v2

@@ G.2 Identifiability of Estimated Bi-factor Structure in Real Data Example
Next, we prove the second part. We denote by $\mathcal{B}_{5}^{*}=\{j_{5},j_{6}\}$, $\mathcal{B}_{6}^{*}=\{j_{7},j_{8}\}$ and $\mathcal{B}_{7}^{*}=\{j_{9},j_{10}\}$. Since $\Lambda$, $\Phi$ and $\Psi$ satisfy Condition 4, there exists three types of possible of bi-factor structure of $\mathcal{B}_{5}$, $\mathcal{B}_{6}$ and $\mathcal{B}_{7}$ and we discuss the three cases one by one. Without loss of generality, we assume $D^{\prime}$ given in the first part equals the identity matrix.
  1. 1.
    None of the bi-factor structures of the variables belonging to $\mathcal{B}_{i}^{*}$, $i=5,6,7$, is correct. Without loss of generality, we assume $\mathcal{B}_{5}=\{j_{5},j_{10}\}$, $\mathcal{B}_{6}=\{j_{6},j_{7}\}$ and $\mathcal{B}_{7}=\{j_{8},j_{9}\}$. In this case, we consider the equation
    $\Sigma[\mathcal{B}_{1}^{*},\mathcal{B}_{5}^{*}]=\Sigma^{*}[\mathcal{B}_{1}^{*},\mathcal{B}_{5}^{*}],$ (G.21)
    which is equivalent to
    $\Lambda[\mathcal{B}^{*}_{1},\{1\}]\lambda_{j_{5},1}+\phi_{2,6}\Lambda[\mathcal{B}^{*}_{1},\{2\}]\lambda_{j_{5},6}=\Lambda^{*}[\mathcal{B}^{*}_{1},\{1\}]\lambda^{*}_{j_{5},1}+\phi^{*}_{2,6}\Lambda^{*}[\mathcal{B}^{*}_{1},\{2\}]\lambda^{*}_{j_{5},6},$ (G.22)
    $\Lambda[\mathcal{B}^{*}_{1},\{1\}]\lambda_{j_{6},1}+\phi_{2,7}\Lambda[\mathcal{B}^{*}_{1},\{2\}]\lambda_{j_{6},7}=\Lambda^{*}[\mathcal{B}^{*}_{1},\{1\}]\lambda^{*}_{j_{6},1}+\phi^{*}_{2,6}\Lambda^{*}[\mathcal{B}^{*}_{1},\{2\}]\lambda^{*}_{j_{6},6}.$
    Since the first part is proved, we have assumed that $\Lambda[\mathcal{B}^{*}_{1},\{1\}]=\Lambda^{*}[\mathcal{B}^{*}_{1},\{1\}]$ and $\Lambda[\mathcal{B}^{*}_{1},\{2\}]=\Lambda^{*}[\mathcal{B}^{*}_{1},\{2\}]$. Noticing that $\Lambda^{*}[\mathcal{B}^{*}_{1},\{1\}]$ and $\Lambda^{*}[\mathcal{B}^{*}_{1},\{2\}]$ are linearly independent, we have $\lambda_{j_{5},1}=\lambda^{*}_{j_{5},1}$ and $\lambda_{j_{6},1}=\lambda^{*}_{j_{6},1}$. Similarly, by considering the equations $\Sigma[\mathcal{B}_{1}^{*},\mathcal{B}_{6}^{*}]=\Sigma^{*}[\mathcal{B}_{1}^{*},\mathcal{B}_{6}^{*}]$ and $\Sigma[\mathcal{B}_{1}^{*},\mathcal{B}_{7}^{*}]=\Sigma^{*}[\mathcal{B}_{1}^{*},\mathcal{B}_{7}^{*}]$, we have $\lambda_{j_{7},1}=\lambda^{*}_{j_{7},1}$, $\lambda_{j_{8},1}=\lambda^{*}_{j_{8},1}$, $\lambda_{j_{9},1}=\lambda^{*}_{j_{9},1}$ and $\lambda_{j_{10},1}=\lambda^{*}_{j_{10},1}$.
    By considering the equation
    $\Sigma[j_{5},j_{6}]=\Sigma^{*}[j_{5},j_{6}],$ (G.23)
    which is equivalent to $\lambda_{j_{5},1}\lambda_{j_{6},1}+\phi_{6,7}\lambda_{j_{5},6}\lambda_{j_{6},7}=\lambda^{*}_{j_{5},1}\lambda^{*}_{j_{6},1}+\lambda^{*}_{j_{5},6}\lambda^{*}_{j_{6},6}$. We further have
    $\phi_{6,7}\lambda_{j_{5},6}\lambda_{j_{6},7}=\lambda^{*}_{j_{5},6}\lambda^{*}_{j_{6},6}.$ (G.24)
    We can similarly have the equations
    $\phi_{6,7}\lambda_{j_{5},6}\lambda_{j_{7},7}$ $=$ $\phi_{6,7}^{*}\lambda^{*}_{j_{5},6}\lambda^{*}_{j_{7},7},$ (G.25)
    $\phi_{6,7}\lambda_{j_{10},6}\lambda_{j_{6},7}$ $=$ $\phi_{6,8}^{*}\lambda^{*}_{j_{10},8}\lambda^{*}_{j_{6},6},$
    $\phi_{6,7}\lambda_{j_{10},6}\lambda_{j_{7},8}$ $=$ $\phi_{7,8}^{*}\lambda^{*}_{j_{10},8}\lambda^{*}_{j_{7},7}.$
    By combining the equations (G.24) and (G.25), we have $\phi^{*}_{6,7}\phi^{*}_{6,8}=\phi^{*}_{7,8}$. In a symmetric manner, we further have $\phi^{*}_{6,8}\phi^{*}_{7,8}=\phi^{*}_{6,7}$ and $\phi^{*}_{6,7}\phi^{*}_{7,8}=\phi^{*}_{6,8}$. Since $\phi^{*}_{6,7},\phi^{*}_{6,8}$ and $\phi^{*}_{7,8}\neq 0$, we have $|\phi^{*}_{6,7}\phi^{*}_{6,8}\phi^{*}_{7,8}|=1$, which leads to $|\phi^{*}_{6,7}|=|\phi^{*}_{6,8}|=|\phi^{*}_{7,8}|=1$ and violates the assumption that $\Phi^{*}$ is positive definite. Thus, this case does not exist.
  2. 2.
    Only one of the bi-factor structures of the variables belonging to $\mathcal{B}_{i}^{*}$, $i=5,6,7$, is correct. Without loss of generality, we assume $\mathcal{B}_{5}=\{j_{5},j_{6}\}$, $\mathcal{B}_{6}=\{j_{7},j_{9}\}$ and $\mathcal{B}_{7}=\{j_{8},j_{10}\}$. By the same argument in the first case, we have $\lambda_{j_{i},1}=\lambda^{*}_{j_{i},1}$ for $i=5,\ldots,10$. Next, consider the equations on the diagonal entries of
    $\Sigma[\mathcal{B}_{6}^{*}\cup\mathcal{B}_{7}^{*},\mathcal{B}_{6}^{*}\cup\mathcal{B}_{7}^{*}]=\Sigma^{*}[\mathcal{B}_{6}^{*}\cup\mathcal{B}_{7}^{*},\mathcal{B}_{6}^{*}\cup\mathcal{B}_{7}^{*}].$ (G.26)
    we have the following 6 equations
    $\phi_{7,8}\lambda_{j_{7},7}\lambda_{j_{8},8}=\lambda^{*}_{j_{7},7}\lambda^{*}_{j_{8},7},$ (G.27)
    $\lambda_{j_{7},7}\lambda_{j_{9},7}=\phi^{*}_{7,8}\lambda^{*}_{j_{7},7}\lambda^{*}_{j_{9},8},$
    $\phi_{7,8}\lambda_{j_{7},7}\lambda_{j_{10},8}=\phi^{*}_{7,8}\lambda^{*}_{j_{7},7}\lambda^{*}_{j_{10},8},$
    $\phi_{7,8}\lambda_{j_{8},8}\lambda_{j_{9},7}=\phi^{*}_{7,8}\lambda^{*}_{j_{8},7}\lambda^{*}_{j_{9},8},$
    $\lambda_{j_{8},8}\lambda_{j_{10},8}=\phi^{*}_{7,8}\lambda^{*}_{j_{8},7}\lambda^{*}_{j_{10},8},$
    $\phi_{7,8}\lambda_{j_{9},7}\lambda_{j_{10},8}=\lambda^{*}_{j_{9},8}\lambda^{*}_{j_{10},8}.$
@@
    According to the first equation above, we have $\phi_{7,8}\neq 0$. By the 6 equations, we also have $(\phi^{*}_{7,8})^{4}\phi_{7,8}^{4}=\phi_{7,8}^{2}$. Then we have $|\phi^{*}_{7,8}|=|\phi_{7,8}|=1$, which violates the assumption that $\Phi^{*}$ is positive definite. Thus, this case does not exist.
  3. 3.
    The bi-factor structure is correct. Without loss of generality, we assume $\mathcal{B}_{5}=\{j_{5},j_{6}\}$, $\mathcal{B}_{6}=\{j_{7},j_{8}\}$ and $\mathcal{B}_{7}=\{j_{9},j_{10}\}$. Similar to the previous argument, we first have $\lambda_{j_{i},1}=\lambda^{*}_{j_{i},1}$ for $i=5,\ldots,10$. Moreover, we have $\phi_{2,6}\lambda_{5,6}=\phi^{*}_{2,6}\lambda^{*}_{5,6}$, $\phi_{2,6}\lambda_{6,6}=\phi^{*}_{2,6}\lambda^{*}_{6,6}$ and $\lambda_{5,6}\lambda_{6,6}=\lambda^{*}_{5,6}\lambda^{*}_{6,6}$. These 3 equations leads to $\phi_{2,6}=\phi^{*}_{2,6}$, $\lambda_{5,6}=\lambda^{*}_{5,6}$ and $\lambda_{6,6}=\lambda^{*}_{6,6}$ or $\phi_{2,6}=-\phi^{*}_{2,6}$, $\lambda_{5,6}=-\lambda^{*}_{5,6}$ and $\lambda_{6,6}=-\lambda^{*}_{6,6}$. With the same argument, the loadings and correlations related with the variables belonging to $\mathcal{B}_{6}^{*}$ and $\mathcal{B}_{7}^{*}$ are also determined up to a sign flip. The check of $\psi_{j}=\psi^{*}_{j}$ for $j\in\mathcal{B}_{i}^{*}$, $i=5,6,7$ are straight forward.